OpenStudy (anonymous):

What is the quotient of the rational expression x^2+5x+6/x-6 divided by x^2-9/2x-12? Also, it needs to be in reduced form.

OpenStudy (anonymous):

Is it : \[\frac{ x^2+5x+6}{x-6}\div \frac{x^2-9}{2x-12}\] ???

OpenStudy (anonymous):

@Noura11 yes I apologize, I didnt know how to create an equation on here.

OpenStudy (anonymous):

All you need is to factor some expressions you have ! Can you factor \[x^2+5x+6\] ?

OpenStudy (anonymous):

I having a hard time understanding what to factor

OpenStudy (anonymous):

It is very simple ! Look ! The expression : \[ x^2+5x+6\] can be factored like this : \[x^2+5x+6=(x+a)(x+b)\] where a and b are two numbers such that : \[a+b=5\text{ and } a\times b=6\] We can easily say that : \[2+3=5\text{ and } 2\times 3=6\] So : a=2 and b=3 (or the inverse !!!) So : \[ x^2+5x+6=(x+2)(x+3)\]

OpenStudy (anonymous):

so the answer would be 2(x+2)/(x-3) correct?

OpenStudy (anonymous):

Yes ! Exactly !

OpenStudy (anonymous):

and that is also in reduced form correct?

OpenStudy (anonymous):

yes